OpenStudy (anonymous):

A 6 kg rock falls from a cliff. Just before it hits the ground its speed is 24 m/s. From what height did it fall?

OpenStudy (shamim):

U hv to use the formula is v^2=u^2+2*g*h

OpenStudy (shamim):

Given u=0 g=9.8m/s^2 v=24m/s h=?

OpenStudy (anonymous):

29.4 m

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

Medal???!!!!!!!

OpenStudy (shamim):

U r most welcome